OpenStudy (anonymous):

Find the value of y when x equals 3 in this equation. -4x + 5y = -2 A. 3 B. 2 C. -2 D. -4

OpenStudy (anonymous):

-4(3)+5y=-2 5y= -2 +12 5y=10 y=2